# Link groups of Kishino knot stacks

### v2

@@ Abstract.
For any virtual link, a class of new links can be defined called stacks, in which copies of the virtual link are placed on top of one another. The resulting virtual link depends only on the virtual isotopy class of the original link, and the fundamental group of such a link may be used to detect whether the link is nontrivial and whether it is nonclassical in some cases. We show that this group is able to distinguish five Kishino knots from the unknot using a stacked pair. However, for two other Kishino knots, the group and quandle of any stack invariant will be free with a number of generators equal to the number of copies in the stack, although the Jones polynomial of the stacks is able to detect their nontriviality.
@@ 1. Introduction
Virtual links were introduced by Kauffman, [1], and given a geometric interpretation by Kuperberg, [2]. The concept of a stack for a virtual link was introduced in [7]. The general idea is to consider the virtual link as a link in a thickened surface, and then to glue copies of the link in the thickened surface below the original link, possibly with some copies having the orientation of the thickening reversed. The virtual isotopy class of the new link depends only on the virtual isotopy class of the original link, and thus the new link allows for constructing invariants for the original.
In particular, the group and quandle of the stack form an invariant for the original link. In [7] it is shown that this can be nontrivial even if the group of the original link is trivial. However, it is remarked that this is not true for a certain Kishino knot, which are not distinguished from the unknot by checking the group.
In this paper we compute the link groups for stacks of seven different Kishino knots, and show that the link group of stacks of two copies of the original knot suffice to distinguish five of these knots from the unknot. This is particularly interesting since such knots have trivial Jones polynomial and are thus rather hard to detect (although this is possible using biquandle counting invariants, [4]. This detection of these three Kishino knots using elementary algebraic methods is our first result.
However, for two other Kishino knots, one of which was studied in [7], the stack of two copies has a link group that is free on two generators. Therefore, the link group (and quandle) fail to distinguish this knot from the unknot. It is therefore natural to ask whether this problem could be alleviated by adding more copies to the stack. We show that this is not the case, and in particular, that a stack of $n$ copies of such a knot will have a link group that is free on $n$ generators. This is the main general result about stack invariants in this paper. For both of these Kishino knots, the Jones polynomials for the stacks are calculated and shown to distinguish them from the unknot.
@@ 2. The Stack Invariant
For a sequence $\langle a_{i}\rangle$ of $+$ and -- symbols, of length $m$, we define the stack invariant of $L$ to be the link $S_{a_{i}}$ to be the link constructed as follows. Let $L_{i}\subset F_{i}\times I$ be a copy of $L$ if $a_{i}=+$ and a copy of $L_{v}$ if $a_{i}=-$. Then identify $F_{i}\times\{0\}$ with $F_{i+1}\times\{1\}$. Finally reparametrize the resulting thickened surface so that the thickening is identified with the unit interval in the obvious way.
@@ 3. Kishino Knot Calculations
There are seven knots which we will consider here, which exhaust the Kishino knots up to horizontal reflection and orientation. Since horizontal reflection and orientation will not affect the link groups, this is sufficient for our purposes, namely, to test the strength of the homotopy invariants of stacks by using the Kishino knots. Note that the vertical reflections may change the link groups of the stacks. However, also note that vertical reflections have related stacks: if $L$ and $L^{\prime}$ are vertical reflections of one another, then $S_{-+}(L)$ is identical to $S_{+-}(L+)$, for example. Nonetheless, we consider the vertical reflections of the Kishino knots separately.
In particular, we will consider $K$, $K_{switch}$, $K_{alt}$, $K_{v}$, $K_{5}$, $K_{6}$, and $K_{7}$. The groups are calculated by checking crossings, and the presentations simplified using Magma, [3]. More generally any (virtual) knot with the same diagram up to switching classical crossings may be termed a Kishino knot.
@@
Figure 2. The knot $K_{switch}$ with the vertical double on the right and the double on the left.
Figure 3. The knot $K_{alt}$ with the vertical double on the right and the double on the left.
Figure 4. The knot $K_{v}$ with the vertical double on the right and the double on the left.
Figure 5. The knot $K_{5}$ with the vertical double on the right and the double on the left.
Figure 6. The knot $K_{6}$ with the vertical double on the right and the double on the left.
Figure 7. The knot $K_{7}$ with the vertical double on the right and the double on the left.
@@
$\langle a,b,c,d,e,f,g,h,i,j,k,l,m|$
$c=b^{a},c=d^{k},k=l^{a},$
$e=d^{a},m=b^{a},m=l^{e},$
$i=h^{a},e=f^{a},i=j^{e},$
$g=f^{k},k=j^{a},g=h^{a}\rangle$
This simplifies to
$\langle a,e,k|ea^{-1}e^{-1}k^{-1}a^{-1}ek^{-1}a=1\rangle$
@@
$\langle a,b,c,d,e,f,g,h,i,j,k,l,m|$
$c=b^{a},d=c^{a},e=d^{m},$
$e=f^{i},f=g^{a},g=h^{a},$
$i=h^{a},j=i^{a},k=j^{g},$
$k=l^{c},l=m^{a},m=b^{a}\rangle$
This simplifies to
$\langle a,c,g|gc^{-1}a^{-1}cacg^{-1}a^{-1}g^{-1}a=1\rangle$
For $S_{+-}(K_{switch})$:
$\langle a,b,c,d,e,f,g,h,i,j,k,l,m|$
$c=b^{a},c=d^{k},k=l^{a},$
[…]
$f=e^{i},f=g^{a},i=h^{a},$
$j=i^{a},g=h^{a},j=k^{g}\rangle$
For $S_{++}(K_{switch})$:
$\langle a,b,c,d,e,f,g,h,i,j,k,l,m|$
$c=b^{a},l=m^{a},k=l^{c},$
$m=b^{a},d=c^{a},e=d^{m},$
$j=i^{e},i=h^{a},e=f^{a},$
$k=j^{a},g=h^{a},f=g^{k}\rangle$
These both simplify to free groups on $a,c$.
@@
$a^{-1}c^{-1}a^{2}c^{-1}a^{-2}c^{-1}a^{2}ca^{-2}caca^{-1}c^{-1}a^{2}c^{-1}a^{-2}ca^{-2}caca^{-1}c^{-1}a^{2}c^{-1}=1\rangle$
@@
$\langle a,b,c,d,e,f,g,h,i,j,k,l,m|$
$c=b^{a},l=k^{c},l=m^{a},$
$d=c^{a},d=e^{m},m=b^{a},$
$i=h^{a},f=e^{i},f=g^{a},$
$j=i^{a},j=k^{g},g=h^{a}\rangle$
This simplifies to
[…]
@@
$\langle a,b,c,d,e,f,g,h,i,j,k,l,m|$
$d=c^{k},k=l^{a},c=b^{a},$
$e=d^{a},l=m^{e},m=b^{a},$
$i=h^{a},e=f^{a},j=i^{e},$
$k=j^{a},f=g^{k},g=h^{a}\rangle$
Which simplifies to
$\langle a,e,k|$
[…]
For $S_{+-}(K_{5})$:
$\langle a,b,c,d,e,f,g,h,i,j,k,l,m|$
$c=b^{a},c=d^{l},l=m^{a},$
$m=b^{a},e=d^{a},e=f^{m},$
$k=j^{a},g=f^{k},g=h^{a},$
$i=j^{a},i=h^{l},l=k^{a}\rangle$
For $S_{++}(K_{5})$:
$\langle a,b,c,d,e,f,g,h,i,j,k,l,m|$
$c=b^{a},j=k^{c},k=l^{a},$
$m=b^{a},l=m^{d},d=c^{a},$
$g=f^{a},h=g^{d},d=e^{a},$
$j=i^{e},e=f^{a},i=h^{a}\rangle$
These both simplify to free groups on two generators.
@@
$h^{-1}a^{2}h^{-1}a^{-2}haha^{-1}h^{-1}a^{2}h^{-1}a^{-1}h^{-1}a^{2}=1\rangle$
@@
$a^{-2}c^{-1}a^{-1}c^{-1}acac^{-1}a^{-1}c^{-1}aca^{-1}c^{-1}a^{-1}caca=1\rangle$
@@
$a^{-1}da^{2}d^{-1}a^{-2}d^{-1}ad^{-1}a^{-1}da^{2}da^{-1}da^{2}=1\rangle$
For all the cases where we have reduced the presentation to a free group with relator, results on one relator groups show that these cannot be written as the free product of the cyclic group on $a$ with another group (see [5]). Hence they are distinct from the link group of a 2-component unlink.
For the case of $K_{switch}$ and $K_{5}$, we may hope that it is possible to obtain a similar result if we use more layers in the stack. Unfortunately, the following result says that this is not possible.
Theorem 1.
Let $K$ be a virtual link with $S_{+-}(K)$ and $S_{++}(K)$ having link groups that are free on $2m$ generators, where $m$ is the number of components of $K$. Then any stack with $n$ layers for $K$ will have a link group that is free on $nm$ generators.
Proof.
Since $S_{+-}(K)$ and $S_{++}(K)$ have link groups that are free on $2m$ generators, we see that the link group for $K$ must be free on $m$ generators. We denote these generators as $a_{11}$ through $a_{1m}$. In calculating the link group for $S_{+-}(K)$ and $S_{++}(K)$, we obtain generators which consist of $a_{1j}$ and a set of $b_{k}$, one $b_{k}$ for each arc of the new lower copy of $K$. The relators will all be of the form $b_{k}^{b_{l}}=b_{r}$ and $b_{k}^{a_{1l}^{\pm 1}}=b_{r}$, and the resulting group reduces to a free group on $2m$ generators.
Now we proceed by induction. Suppose that our result is established for stacks of $n-1$ layers. Let the generators of the free group for this stack be $a_{ij}$ and let $A_{j}=a_{1j}^{\epsilon_{1}}a_{2j}^{\epsilon_{2}}...a_{(n-1),j}^{\epsilon_{n-1}}$, where $\epsilon_{i}$ is 1 if the $i$-th level of the stack is +, and -1 if it is -. Adding an $n$-th copy of $K$ to the bottom of the stack gives us a link with a group generated by the $a_{ij}$ and a set of $b_{k}$, one $b_{k}$ for each arc of the lowest copy of $K$, together with generators come from the arcs of the new copy that go between the strands of the old copies and can be immediately eliminated, see Fig. 8. The relators will all be of the form $b_{k}^{b_{l}}=b_{r}$ and $b_{k}^{A_{l}^{\pm 1}}=b_{r}$. These relators are identical to those obtained in our first step if we replace $a_{1l}$ by $A_{l}$, and hence, using the same Tietze moves, we can reduce the link group to a free group with $nm$ generators.
Figure 8. Above, the crossing gives a relation for a double stack of $b_{k}^{a_{1l}=b_{r}}$. Below we obtain for the $n$-stack $b_{k}^{A_{l}}=b_{r}$. The thinner arcs have generators that can be immediately removed from the presentation.
Remark 2.
The quandle will also fail to distinguish links such as $K_{switch}$ from unlinks, because the quandle of a link is determined by the group and the peripheral structure.
It is worth noting that the biquandle can distinguish $K_{switch}$ from the unknot, [4].
Calculating the Jones polynomial of $S_{++}(K_{switch})$ yields $-(1+t)(1-3t+3t^{2}-t^{3}+3t^{4}-3t^{5}+t^{6})/t^{7/2}$, [6]. The PD code was found and sent to Adam Sikora, who used it to calculate this polynomial in Mathematica. The Jones polynomial of $S_{+-}(K_{switch})$ was identical to that of the two-component unlink, however.
@@
We conclude, therefore, that two-layer stacks are able to detect the nontriviality of all Kishino knots. For five of the cases (and their horizontal reflections) it is possible to accomplish this using homotopy invariants, while the remaining two cases require us to use quantum invariants.
@@
Question 4.
Is it coincidental that for both $K_{switch}$ and $K_{5}$, both the double and vertical double had trivial group invariants? Does there exist a virtual link $L$ such that the group of the double (or vertical double) is trivial, but the group of the vertical double (or double) is nontrivial?

### v3

# On the strength of invariants from stacks of virtual knots

@@ Abstract.
For any virtual link, a class of new links can be defined called stacks, in which copies of the virtual link are placed on top of one another. The resulting virtual link depends only on the virtual isotopy class of the original link, and the fundamental group of such a link may be used to detect whether the link is nontrivial and whether it is nonclassical in some cases. We show that the groups constructed using this method are sufficient to distinguish all the Kishino knots from the unknot and from one another, as well as calculating their Jones polynomials.
Key words and phrases:
knots, virtual knots, algebraic invariants
1991 Mathematics Subject Classification:
57K12
@@ 1. Introduction
Virtual links were introduced by Kauffman [1], and given a geometric interpretation by Kuperberg [3]. The concept of a stack for a virtual link was introduced in [8]. The general idea is to consider the virtual link as a link in a thickened surface, and then to glue copies of the link in the thickened surface below the original link, possibly with some copies having the orientation of the thickening reversed. The virtual isotopy class of the new link depends only on the virtual isotopy class of the original link, and thus the new link allows for constructing invariants for the original.
In particular, the group and quandle of the stack form an invariant for the original link. In [8] it is shown that this can be nontrivial even if the group of the original link is trivial. However, it is remarked that this is not true for two layer stacks of a certain Kishino knot (herein named $K_{switch}$) which is not distinguished from the unknot by the groups of these two layer stacks. The motivation for introducing stacks in [8] was to attempt to construct topological homotopy invariants that would be more powerful than the usual group and quandle, and as such the question was posed therein whether it was possible to use homotopy invariants to distinguish $K_{switch}$ from the unknot by using more layers in the stack. Our main result is that this is possible using a three layer stack. To our knowledge this is the first use of a homotopy invariant to detect the nontriviality of this knot, although this is possible using biquandle counting invariants [5] or colored Jones polynomials. In addition the Jones polynomials of the stacks are calculated, which distinguish all seven cases using two layer stacks.
@@ 2. The Stack Invariant
For a sequence $\langle a_{i}\rangle.$ of ++ and - symbols, of length $m$, we define the stack invariant of $L$ to be the link $S_{a_{i}}$ to be the link constructed as follows. Let $L_{i}\subset F_{i}\times I$ be a copy of $L$ if $a_{i}=+$ and a copy of $L_{v}$ if $a_{i}=$-. Then identify $F_{i}\times\{0\}$ with $F_{i+1}\times\{1\}$. Finally reparametrize the resulting thickened surface so that the thickening is identified with the unit interval in the obvious way. We refer to $L_{i}$ as the $i$-th layer of the stack, and refer to $m$ as the number of layers.
@@ 3. Kishino Knot Calculations
There are seven knots which we will consider here, which exhaust the Kishino knots up to horizontal reflection and orientation. Since horizontal reflection and orientation will not affect the link groups of the stacks (or the nontriviality of the Jones polynomials considered later), this is sufficient for our purposes, namely, to test the strength of the homotopy invariants of stacks by using the Kishino knots. Note that the vertical reflections may change the link groups of the stacks. However, also note that vertical reflections have related stacks: if $L$ and $L^{\prime}$ are vertical reflections of one another, then $S_{-+}(L)$ is identical to $S_{+-}(L^{\prime})$, for example. Nonetheless, we consider the vertical reflections of the Kishino knots separately, since we are interested in whether invariants from stacks can distinguish them.
In particular, we will consider $K$, $K_{switch}$, $K_{alt}$, $K_{v}$, $K_{5}$, $K_{6}$, and $K_{7}$. The groups are calculated by checking crossings, and the presentations simplified using Magma [4], which was also used to count inequivalent group epimorphisms. Note that all these knots have cyclic knot groups except $K_{v}$, which has a group given by
$\langle a,d|dad^{-1}ad^{-1}a^{-1}=1\rangle.$
Note that this alone shows that not only $K_{v}$ but also $K$ is nontrivial and both are distinct from the other five: $K$ and $K_{v}$ are vertical reflections of one another, but none of the others have nontrivial groups for themselves or their vertical reflections.
@@
Figure 2. The other six Kishino knots considered in this paper.
@@
$\langle a,e,k|ea^{-1}e^{-1}k^{-1}a^{-1}ek^{-1}a=1\rangle.$
This has 75 inequivalent epimorphisms to $S_{5}$.
@@
$\langle a,c,g|gc^{-1}a^{-1}cacg^{-1}a^{-1}g^{-1}a=1\rangle.$
This has 75 inequivalent epimorphisms to $S_{5}$.
The groups for $S_{+-}(K_{switch})$ and $S_{++}(K_{switch})$ both simplify to free groups on two generators.
@@
$a^{-1}c^{-1}a^{2}c^{-1}a^{-2}c^{-1}a^{2}ca^{-2}caca^{-1}c^{-1}a^{2}c^{-1}a^{-2}ca^{-2}caca^{-1}c^{-1}a^{2}c^{-1}=1\rangle.$
This has 12 inequivalent epimorphisms to $S_{5}$.
@@
$\langle B,D,c,g|$
$DB^{-1}DB^{-1}D^{-1}B=1,$
$B^{-2}DcD^{-1}B^{2}c^{-1}gB^{-2}Dg^{-1}D^{-1}B^{2}g^{-1}c=1,$
$B^{-1}Dg^{-1}cB^{-1}D^{-1}BcDB^{-1}Dc^{-1}gD^{-1}BD^{-1}g^{-1}D=1\rangle.$
This has 241 inequivalent epimorphisms to $S_{5}$.
@@
$\langle B,D,e,k|$
$BD^{-1}BD^{-1}B^{-1}D=1,$
$BeB^{-1}k^{-1}eBk^{-1}B^{-1}e^{-1}k=1,$
$DeD^{-1}k^{-1}D^{-1}BkBe^{-1}B^{-1}k^{-1}B^{-1}Dk=1\rangle.$
This has 87 inequivalent epimorphisms to $S_{5}$.
The groups of $S_{+-}(K_{5})$ and $S_{++}(K_{5})$ both simplify to free groups on two generators.
@@
$h^{-1}a^{2}h^{-1}a^{-2}haha^{-1}h^{-1}a^{2}h^{-1}a^{-1}h^{-1}a^{2}=1\rangle.$
This has 14 inequivalent epimorphisms to $S_{5}$.
@@
$a^{-2}c^{-1}a^{-1}c^{-1}acac^{-1}a^{-1}c^{-1}aca^{-1}c^{-1}a^{-1}caca=1\rangle.$
This has 15 inequivalent epimorphisms to $S_{5}$.
@@
$a^{-1}da^{2}d^{-1}a^{-2}d^{-1}ad^{-1}a^{-1}da^{2}da^{-1}da^{2}=1\rangle.$
This has 14 inequivalent epimorphisms to $S_{5}$.
For all the cases where we have reduced the presentation to a free group with one relator, results on one relator groups show that these cannot be written as the free product of the cyclic group on $a$ with another group (see [6]). Hence they are distinct from the link group of a 2-component unlink. Furthermore, by counting inequivalent epimorphisms to $S_{5}$, it may be seen that the nontrivial groups for the vertical doubles and the doubles are all distinct from one another (and are also distinct from the free group on two generators).
$S_{+-}(K)$ 75
$S_{++}(K)$ 75
$S_{+-}(K_{switch})$ free
$S_{++}(K_{switch})$ free
$S_{+-}(K_{alt})$ 12
$S_{++}(K_{alt})$ 12
$S_{+-}(K_{v})$ 241
$S_{++}(K_{v})$ 87
$S_{+-}(K_{5})$ free
$S_{++}(K_{5})$ free
$S_{+-}(K_{6})$ 14
$S_{++}(K_{6})$ 15
$S_{+-}(K_{7})$ 15
$S_{++}(K_{7})$ 14
Table 1. A table showing the inequivalent epimorphisms from the groups to $S_{5}$. Groups marked ”free” are free on two generators.
For the case of $K_{switch}$ and $K_{5}$, we can go to a three layer stack to distinguish these from the unknot (and from one another). This positively answers a question in [8] about whether $K_{switch}$ can be detected using homotopy groups from stacks.
The group for $S_{+++}(K_{switch})$:
$\langle a,c,C|$
$a^{2}c^{-1}a^{-1}caca^{-1}C^{-1}c^{-1}a^{-1}CacCac^{-1}a^{-1}c^{-1}aca^{-2}C^{-1}c^{-1}a^{-1}C^{-1}$
$acCc^{-1}a^{-1}c^{-1}acac^{-1}a^{-1}c^{-1}aca^{-1}c^{-1}a^{-1}cacC^{-1}c^{-1}a^{-1}CacC=1\rangle.$
This has 94 inequivalent epimorphisms to $S_{5}$.
The group for $S_{+++}(K_{5})$ has 928 inequivalent epimorphisms to $S_{5}$.
$K_{1}$, $K_{switch}$, and $K_{v}$ all have trivial Jones polynomials, and require the use of colored Jones to detect, although the other four can be detected by standard Jones polynomials.
$K_{alt}$: $t^{-5}-2t^{-4}-2t^{-7/2}+t^{-3}+2t^{-5/2}+t^{-2}$.
$K_{5}$: $3-t^{-3/2}-t^{-1}+t^{-1/2}+t^{1/2}-t-t^{3/2}$.
$K_{6}$: $t^{-5/2}+t^{-3/2}+t^{-1}$.
$K_{7}$: $t^{-5/2}+t^{-3/2}+t^{-1}$.
[…]
@@
For completeness we list the Jones polynomials of the other five examples considered.
For $S_{++}(K)$, $(-1+2t-2t^{3}-2t^{4}+2t^{6}-t^{7})/t^{7/2}$.
For $S_{+-}(K)$, $(-1-t)/t^{1/2}$. For $S_{++}(K_{v})$, $(-1+2t-2t^{3}-2t^{4}+2t^{6}-t^{7})/t^{7/2}$.
For $S_{+-}(K_{v})$, $(-4+8t-5t^{3}-5t^{4}+8t^{6}-4t^{7})/t^{7/2}$. For $S_{++}(K_{alt})$, $(-1+7t-18t^{2}+20t^{3}-15t^{4}+21t^{5}-19t^{6}+7t^{7}-4t^{8})/t^{21/2}$.
For $S_{+-}(K_{alt})$, $(-1+5t^{2}+t^{3}-6t^{4}-6t^{5}+t^{6}+5t^{7}-t^{9})/t^{9/2}$.
For $S_{++}(K_{6})$, $(-3+9t-7t^{2}+4t^{3}-9t^{4}+6t^{5}-2t^{6})/t^{11/2}$. For $S_{+-}(K_{6})$, $(1-2t^{2}-2t^{3}+t^{5})/t^{5/2}$.
For $S_{++}(K_{7})$, $(-3+9t-7t^{2}+4t^{3}-9t^{4}+6t^{5}-2t^{6})/t^{11/2}$.
For $S_{+-}(K_{7})$, $(-2+5t-4t^{3}-4t^{4}+5t^{6}-2t^{7})/t^{7/2}$.
We see therefore that the seven Kishino knots are distinguished from each other, and from the unknot, by the Jones polynomials of the two layer stacks $S_{+-}$ and $S_{++}$.
Question 1.
Note that $K_{6}$ and $K_{7}$ have the same Jones polynomial.
Question 2.
Is it coincidental that for both $K_{switch}$ and $K_{5}$, both the double and vertical double have trivial group invariants, whereas for the other five, both the double and vertical double have nontrivial group invariants? Does there exist a virtual link $L$ such that the group of the double (or vertical double) is indistinguishable from that of the unlink, but the group of the vertical double (or double) is distinguishing?
The particular interest of this study lies in the fact that, while all of the knots studied here can be classified by means such as colored Jones polynomials or biquandle counting invariants, this is the first time to our knowledge that a classical algebraic topology invariant has been used to classify them. This leads us to a last question:
@@
Are the groups of stacks as strong (or stronger) invariants as the biquandle or the (colored) Jones polynomial invariants?